# KineSoft: Learning Proprioceptive Manipulation Policies with Soft Robot Hands

### v1

@@ Abstract
Underactuated soft robot hands offer inherent safety and adaptability advantages over rigid systems, but developing dexterous manipulation skills remains challenging. While imitation learning shows promise for complex manipulation tasks, traditional approaches struggle with soft systems due to demonstration collection challenges and ineffective state representations. We present KineSoft, a framework enabling direct kinesthetic teaching of soft robotic hands by leveraging their natural compliance as a skill teaching advantage rather than only as a control challenge. KineSoft makes two key contributions: (1) an internal strain sensing array providing occlusion-free proprioceptive shape estimation, and (2) a shape-based imitation learning framework that uses proprioceptive feedback with a low-level shape-conditioned controller to ground diffusion-based policies. This enables human demonstrators to physically guide the robot while the system learns to associate proprioceptive patterns with successful manipulation strategies. We validate KineSoft through physical experiments, demonstrating superior shape estimation accuracy compared to baseline methods, precise shape-trajectory tracking, and higher task success rates compared to baseline imitation learning approaches. KineSoft’s results demonstrate that embracing the inherent properties of soft robots leads to intuitive and robust dexterous manipulation capabilities. Videos and code will be available at kinesoft-policy.github.io.
I Introduction
@@
Underactuated soft robotic hands offer significant advantages over rigid counterparts, including inherent safety through material compliance [49] and robust adaptability to uncertain object geometries [6, 14, 43]. These properties make them particularly well-suited for applications requiring close human-robot interaction, such as assistive robotics, and collaborative manufacturing [49, 11]. However, imparting dexterous in-hand manipulation skills to soft hands remains challenging. Traditional methods for soft robot manipulation often rely on hand-crafted primitives [6, 43, 1] that necessitate expert operators and limit system adaptability.
@@
Recent advances in imitation learning, particularly frameworks like diffusion policy, have shown promise in teaching complex manipulation skills [8, 22, 15]. These approaches have been successfully applied to various tasks, from long-horizon mobile manipulation with rigid grippers [12] to deformable object manipulation with simple end-effectors [47]. Unlike reinforcement learning methods that depend on carefully crafted reward functions and simulation environments [26, 36], imitation learning requires only demonstration trajectories of successful task execution. However, collecting such demonstrations for soft hands presents challenges: traditional demonstration collection methods generally do not apply to underactuated soft systems with virtually infinite degrees of freedom. Conventional teleoperation interfaces [27], designed for rigid anthropomorphic hands, fail to capture the unique capabilities and constraints of underactuated soft end-effectors, which often lack an intuitive mapping to rigid human hand joints. Additionally, standard robot state representations for rigid robots in imitation learning frameworks [50], such as rigid transformation poses, struggle to provide meaningful state information when applied to continuously deforming structures. Despite recent advancements in expressive state representation learning for soft robots [37, 46, 48], these have not yet been applied to skill learning frameworks for in-hand manipulation. These limitations have restricted the application of imitation learning to soft robotic manipulation.
@@
We present KineSoft, a hierarchical framework that enables direct kinesthetic teaching of soft robotic hands. Our key insight is that soft robots’ natural compliance is an advantage for teaching rather than just a control challenge. This compliance allows human demonstrators to physically guide the soft robot fingers through desired movements, enabling intuitive demonstration collection that naturally accounts for the system’s mechanical properties without fighting against kinematic constraints. As shown in Figure LABEL:fig:front, KineSoft is composed of three key components. First, the proprioceptive system achieves state-of-the-art shape estimation using internal strain sensing arrays and a model trained on large simulated data of the robot’s high-dimensional configurations. These sensors provide rich proprioceptive feedback while preserving the hand’s natural compliance, allowing KineSoft to capture detailed information about the hand’s deformation state during manipulation tasks in real time. Next, we train an imitation policy on these shape trajectories and use it to generate deformation trajectories during rollout. Finally, KineSoft’s low-level shape-conditioned controller then tracks these desired shapes. Experiments demonstrate that KineSoft achieves accurate shape state estimation, precise trajectory tracking through the shape-conditioned controller, and superior performance in learned manipulation tasks through these shape-based representations.
In summary, this paper provides the following contributions:
  • KineSoft, a framework for learning from kinesthetic demonstrations for soft robot hands that enables dexterous in-hand manipulation,
@@
  • A state-of-the-art proprioceptive shape estimation approach using strain sensing integrated with soft robot hands that enables precise tracking of finger deformations during contact-rich tasks,
  • Shape-conditioned controller for tracking the generated deformation trajectories and performing dexterous manipulation tasks, and
  • Simulated dataset and trained model for state estimation and control, which we demonstrate can be readily deployed to open-source soft robot hands, such as the MOE platform [49].
Figure 1: Simulated Workspace of the Soft Finger. Top: The soft finger’s workspace is visualized through sampled tip positions that form a semi-hemispherical shape, demonstrating its wide range of motion. Different colors represent different finger configurations. Bottom: example simulated sensor strain and corresponding mesh vertex displacements.
II Related Work
II-A Learning for In-hand Dexterity
@@
Imitation learning has emerged as a promising alternative for reducing the reliance on explicit physics simulation environments, enabling robots to acquire manipulation skills efficiently with real-world data [16, 8, 50]. In-hand manipulation skills through imitation learning are typically achievable with anthropomorphic hands that provide an intuitive mapping between human hand and robot poses [3, 4, 39]. However, despite inherent benefits in safety and dexterity through compliance, which soft robot hands naturally provide, soft robots were subject to some unique challenges—particularly due to the absence of reliable proprioceptive feedback [40] and thus the lack of practical frameworks for collecting demonstrations for soft robots. To the best of our knowledge, KineSoft is the first framework that effectively leverages passive compliance of the soft robots to collect demonstrations and enables soft robots to acquire dexterous in-hand manipulation skills.
Figure 2: Robot setup. A three-fingered MOE hand is mounted on a 6 degree-of-freedom robot arm. A RGB-D camera is mounted on the wrist of the arm to provide object-interaction observations. During domain alignment and shape-tracking evaluation phases, we project random color patterns on the MOE fingers to add visual features for the RGB-D camera.
II-B Soft Robot Dexterity
@@
Soft robot hands excel in grasping and manipulation tasks due to their material compliance, which allows passive adaptation to diverse object geometries [28, 53]. This compliance facilitates robust grasping and safe interactions with humans and delicate objects by distributing contact forces across large areas [1, 18, 49]. Recent advancements in soft robotics have aimed to enhance dexterity through innovative actuator designs, material improvements, and bio-inspired morphologies [25, 11]. Despite these strides, controlling soft hands remains a significant challenge due to their complex dynamics and the high-dimensional nature of their deformation spaces [44]. Consequently, learning-based approaches for soft robotic dexterity have primarily focused on grasping [13], while the development of dexterous in-hand manipulation skills has been hindered by the lack of intuitive demonstration methods and reliable proprioceptive feedback [40]. The unique advantages of soft robot hands in robust in-hand manipulation stem from their lack of rigid skeletal structures [23]. However, this absence also introduces challenges, as soft robot kinematics differ significantly from human hand motions. The KineSoft framework bridges these gaps by integrating novel and accurate shape estimation methods with learned imitation policies, enabling efficient skill acquisition for dexterous manipulation with soft robots.
II-C Soft Robot State Representation
@@
Proprioceptive shape sensing is critical for enabling dexterous control in soft robots, particularly for accurate shape tracking and feedback-driven control [52]. Existing works often utilize low degree-of-freedom shape representations such as constant curvature models [32, 45] or bending angles [35], which fail to capture the full richness of soft robot deformation states. Toward capturing these complex deformation behaviors of soft robot manipulators, recent approaches based on mechanics models have employed Cosserat rod models and high-dimensional Frenet–Serret frames, associated with the continuum cross sections [20]. However, these approaches are computationally costly to update and to preserve hard constraints [20, 19]. Recent learning-based models have introduced more expressive representations of soft robot states using point clouds [37, 46] and meshes [48, 33]. However, these learned representations have not been connected to policy learning for dexterous manipulation tasks. Addressing this gap, KineSoft proposes a novel framework that leverages proprioceptive sensing and learned shape representations, based on vertex displacement fields over meshes, to facilitate dexterous in-hand manipulation skill learning in soft robots.
Figure 3: Proprioception Network Architecture. Sensors are embedded compactly between the tendons. KineSoft’s proprioception learns a mapping from the observed embedded sensors’ signals to the deformation and displacement of the mesh vertices conditioned on the initial points of the mesh.
III Problem Statement
@@
We consider the problem of dexterous in-hand manipulation with a soft robotic hand. Let $\mathbf{M}$ represent the true underlying continuous deformation state space of the hand. The hand is actuated through a set of control inputs $\mathbf{U}$, with the relationship between actuation and deformation being governed by the hand’s material properties and mechanical design. The problem of learning dexterous manipulation skills consists of key challenges: i) we need an observable surrogate $\mathbf{S}$ of the true deformation state space, ii) we need a reliable mapping between the surrogate space $\mathbf{S}$ and the true deformation state space $\mathbf{M}$, and iii) we need a policy $\pi$, which maps the current state to appropriate control actions: $\pi:\mathbf{M}\times\mathcal{O}\mapsto\mathbf{U}$, where $\mathcal{O}$ represents additional task-relevant observations. This presents unique challenges for soft robots compared to rigid systems: the continuous deformation space $\mathbf{M}$ is theoretically infinite-dimensional as a continuum, demonstrations must account for the robot’s inherent compliance, and the mapping between actuation and deformation is highly nonlinear or difficult to simulate in its entirety [20]. The objective is to develop a framework that can effectively learn and execute manipulation skills while embracing these fundamental characteristics of soft robots.
IV Method
IV-A MOE Soft Robot End-effector
@@
In this work, we leverage the MOE soft robot platform, introduced by Yoo et al. [49], which comprises modular finger units that each operate independently. Each finger is actuated by two servo motors, which apply tension to four tendons, as shown in Figure 3. This modular design provides flexibility, allowing the fingers to be rearranged into various configurations to suit specific task requirements. In this study, we examine a three-finger variant of MOE, inspired by research on object controllability using three-fingered rigid end-effectors [21].
@@
Building on the original MOE finger design, we present an enhancement by embedding low-cost conductive elastic rubber directly into the silicone elastomer body of each finger. These sensors measure deformation by varying their electrical resistance as they stretch, providing real-time proprioceptive feedback. Each finger incorporates four of these sensors, compactly positioned between the tendons. The sensors are connected to a data acquisition (DAQ) circuit and board capable of recording resistance readings at approximately 400 Hz. By seamlessly integrating the sensors into the elastomer during fabrication, we developed a fully deformable and sensorized finger body. The combined state spaces from all twelve (12) strain sensors provides our estimate $\mathbf{S}$ of the soft robot hand’s true deformation state space $\mathbf{M}$ (Section III).
In contrast to prior works [33, 24], which relied on custom sensor fabrication or specialized materials, our proposed modifications to the MOE design leverage off-the-shelf conductive rubber, eliminating the need for complex manufacturing processes, enabling straightforward integration into robotic systems, and enhancing accessibility and scalability across a wide range of applications. Furthermore, we present a shape estimation model trained on large simulated soft robot dataset, which can be deployed on sensorized MOE hands through autonomous domain alignment as discussed in later sections.
Figure 4: Domain alignment results. We iteratively update the coefficients to align the real-world sensor signal distribution to the simulated sensor signals. After 200 iterations, we converged to a solution.
IV-B Shape Estimation Model

### v2

@@ Abstract
Underactuated soft robot hands offer inherent safety and adaptability advantages over rigid systems. While imitation learning shows promise for acquiring complex dexterous manipulation skills, adapting existing methods to soft robots presents unique challenges in state representation and data collection. We propose KineSoft, a framework for direct kinesthetic teaching of soft robotic hands that leverages their natural compliance as a skill teaching advantage rather than only as a control challenge. With KineSoft, human demonstrators physically guide the robot while the system learns to associate proprioceptive patterns with successful manipulation strategies. KineSoft makes three key contributions: (1) a shape-based imitation learning framework that uses proprioceptive feedback to ground diffusion-based policies (2) a low-level shape-conditioned controller that enables precise tracking of desired shape trajectories, (3) a sim-to-real learning approach to soft robot mesh shape sensing with an internal strain-sensing array. In physical experiments, we demonstrate the superiority of KineSoft over the strain-based policy baseline in six in-hand manipulation tasks involving both rigid and deformable objects. KineSoft’s results suggest that embracing the inherent properties of soft robots leads to intuitive and robust dexterous manipulation capabilities. Videos and code will be available upon final decision at kinesoft-policy.github.io.
Keywords: Soft Robots, Proprioceptive Estimation, Imitation Learning
Figure 1: KineSoft is a framework for learning from kinesthetic demonstration, enabling free-shaped soft end-effectors to perform dexterous manipulation. Three key components are: 1) a proprioceptive model for high-fidelity shape estimation, 2) diffusion-based imitation learning for predicting the changes in shape and end-effector poses, and 3) a shape-conditioned controller that allows the soft hand to track given shape trajectories.
1 Introduction
@@
Underactuated soft robot hands offer key advantages over rigid counterparts, including inherent safety through material compliance [1, 2] and robust adaptability to uncertain object geometries [3, 4, 5]. These properties make them well-suited for applications requiring contact-rich and reliable interactions such as assistive care, delicate object manipulation, fruit picking, and collaborative manufacturing [1, 6]. However, imparting dexterous in-hand manipulation skills to soft hands remains challenging. Existing methods for soft robot manipulation often rely on hand-crafted primitives [3, 5, 7, 8] that necessitate expert operators and limit system adaptability.
@@
Recent advances in imitation learning, particularly frameworks like diffusion policy, have shown promise in teaching complex manipulation skills [9, 10, 11]. These approaches have been successfully applied to various tasks, from long-horizon mobile manipulation with rigid grippers [12] to deformable object manipulation with simple end-effectors [13]. Unlike reinforcement learning methods that depend on carefully crafted reward functions and simulation environments [14, 15], imitation learning requires only demonstration trajectories of successful task execution. However, collecting such demonstrations for soft robots presents challenges: demonstration collection methods for rigid articulated robots generally do not apply to underactuated soft robots and their virtually infinite degrees of freedom. Teleoperation interfaces [16] designed for rigid anthropomorphic hands fail to capture the unique capabilities and constraints of underactuated soft end-effectors. Soft robots often lack a direct mapping to rigid human hand joints. Additionally, standard robot state representations for rigid robots in imitation learning frameworks [17], such as rigid transformation poses, are ill-suited to provide meaningful state information for continuously deforming structures. Despite recent advances in expressive state representation learning for soft robots [18, 19, 20], these have not yet been applied to skill learning frameworks for in-hand manipulation. These limitations have restricted the application of imitation learning to soft robotic manipulation.
@@
We propose KineSoft, a hierarchical framework to enable direct kinesthetic teaching of soft robot hands skills. The key insight is that soft robots’ natural compliance provides an advantage for teaching rather than just a control challenge. Unlike rigid robots, whose joints and linkages often resist human guidance due to strict kinematic constraints, soft hands can be easily deformed by a human demonstrator into desired poses. This allows demonstrations to be collected by physically manipulating the hand without violating mechanical limits or producing infeasible configurations. KineSoft has three key components (Figure 1). (1) The proprioceptive system achieves state-of-the-art shape estimation using internal strain-sensing arrays and a model trained on large simulated data of the robot’s high-dimensional configurations. These sensors provide rich proprioceptive feedback while preserving the hand’s natural compliance, allowing KineSoft to capture detailed information about the hand’s deformation state during manipulation tasks in real time. (2) We train an imitation policy on these shape trajectories and use it to generate deformation trajectories during rollout. (3) KineSoft’s low-level shape-conditioned controller then tracks these desired shapes. Experiments demonstrate that KineSoft achieves accurate shape state estimation, mesh-based trajectory tracking through the shape-conditioned controller, and high performance in learned manipulation skills through these shape-based representations.
@@
In summary, this paper contributes: i) KineSoft, a framework for learning from kinesthetic demonstrations for soft robot hands that enables dexterous in-hand manipulation, ii) a state-of-the-art proprioceptive shape estimation approach using strain sensing integrated with soft robot hands that enables precise tracking of finger deformations during contact-rich tasks, iii) shape-conditioned controller for tracking the generated deformation trajectories and performing dexterous manipulation tasks, and iv) simulated dataset and trained model for state estimation and control, which we demonstrate can be readily deployed to open-source soft robot hands, such as the MOE platform [1].
2 Related Work
Learning for in-hand dexterity.
@@ 2 Related Work
Imitation learning has emerged as a promising alternative for reducing the reliance on explicit physics simulation environments, enabling robots to acquire manipulation skills efficiently with real-world data [26, 9, 17]. In-hand manipulation skills through imitation learning are typically achievable with anthropomorphic hands that provide a direct mapping between human hand and robot poses [27, 28, 29]. However, despite the inherent benefits in safety and dexterity through compliance of soft robot hands, they face unique challenges due to the absence of reliable proprioceptive feedback [30] and thus there is a lack of practical frameworks for collecting demonstrations for soft robots. To the best of our knowledge, KineSoft is the first framework that effectively leverages passive compliance of soft robots to collect demonstrations and enables soft robots to acquire dexterous in-hand manipulation skills.
Soft robot dexterity.
@@
Soft robot hands excel in grasping and manipulation tasks due to their material compliance, allowing for passive adaptation to diverse object geometries [31, 32]. This compliance facilitates robust grasping and safe interactions with humans and delicate objects by distributing contact forces [7, 33, 1]. Recent advances in soft robots have aimed to improve dexterity through innovative actuator designs, material improvements, and bioinspired morphologies [34, 6]. Despite these strides, controlling soft hands remains a significant challenge due to their complex dynamics and the high dimensionality of their state spaces [35]. Consequently, learning-based approaches for soft robot dexterity have primarily focused on grasping [36], while the development of dexterous in-hand manipulation skills has been hindered by the lack of available demonstration methods and reliable proprioceptive feedback [30]. The unique advantages of soft robot hands in robust in-hand manipulation stem from their lack of rigid skeletal structures [37]. However, this absence also introduces challenges, as soft robot kinematics differ significantly from human hand motions. The KineSoft framework bridges these gaps by integrating novel and accurate shape estimation methods with learned imitation policies, enabling efficient skill acquisition for dexterous manipulation with soft robots.
Soft robot state representation.
@@
Proprioceptive shape sensing is critical for enabling robust control in soft robots, particularly for accurate shape tracking and feedback-driven control [38, 39]. Existing works often use low degree-of-freedom shape representations such as constant curvature models [40, 41] or bending angles [42], which fail to capture the full richness of soft robot deformation states. Toward capturing these complex deformation behaviors of soft robot manipulators, recent approaches based on mechanics models have employed Cosserat rod models and high-dimensional Frenet-Serret frames, associated with the continuum cross sections [43]. However, updating and preserving hard constraints using these approaches is computationally expensive [43, 44]. Recent learning-based models have introduced more expressive representations of soft robot states using point clouds [18, 19] and meshes [20, 45]. However, these learned representations have not been connected to policy learning for dexterous manipulation tasks. Addressing this gap, KineSoft proposes a novel framework that leverages proprioceptive sensing and learned shape representations, based on vertex displacement fields over meshes, to facilitate dexterous manipulation skill learning in soft robots.
Figure 2: Proprioception network. A: Network architecture for mesh shape estimation of the soft fingers. B: Results of domain alignment iterations where the loss converged after 200 iterations.
3 Problem Statement
We consider the problem of learning dexterous manipulation skills with a soft robot hand equipped with internal proprioceptive sensing. Let $\mathbf{M}$ denote the true continuous deformation state of the hand, which is governed by the robot’s mechanical structure and actuated through control inputs $\mathbf{U}$. Since $\mathbf{M}$ is unobservable and high- high-dimensional [46] in practice, we assume access to an internal sensor measurement space $\mathbf{S}$ (e.g., from strain sensors) that provides indirect information about the deformation.
The objective is to learn a policy $\pi:\mathbf{M}\times\mathcal{O}\rightarrow\mathbf{U}$ that maps from the current deformation state and additional task-relevant observations $\mathcal{O}$ (e.g., visual input) to a control action in $\mathbf{U}$. This setting poses three key challenges: i) Estimating the unobservable deformation state $\mathbf{M}$ from the sensor measurement space $\mathbf{S}$; ii) Learning a reliable mapping from $\mathbf{S}$ to $\mathbf{M}$ for downstream use in control; iii) Learning a control policy $\pi$ that uses $\mathbf{M}$ and $\mathcal{O}$ to generate actions, where $\mathcal{O}$ may include exteroceptive modalities (e.g., vision) that are complementary to the proprioceptive signals in $\mathbf{S}$.
@@ 3 Problem Statement
This presents unique challenges for soft robots compared to rigid systems: the continuous deformation space $\mathbf{M}$ is theoretically infinite-dimensional as a continuum, demonstrations must account for the robot’s inherent compliance, and the mapping between actuation and deformation is nonlinear or difficult to simulate in its entirety [43]. The objective is to develop a framework that can effectively learn and execute manipulation skills while embracing these fundamental characteristics.
4 Method
KineSoft enables soft robot dexterous manipulation through a hierarchical approach that bridges the gap between kinesthetic teaching and autonomous execution. Our framework comprises five integrated components: a sensorized MOE-based soft hand with embedded strain sensors, a high-fidelity mesh-based shape estimation model, a shape-conditioned controller for precise trajectory tracking, a diffusion-based shape imitation policy, and a low-level controller that tracks desired shapes. By representing the soft hand’s state through geometric mesh deformations rather than raw sensor values, KineSoft leverages the natural compliance of soft robots for dexterous skill learning while maintaining the precision necessary for complex manipulation tasks.
4.1 MOE Soft Robot End-effector
@@
We leverage the multifinger omnidirectional end-effector (MOE) soft robot platform [1], which comprises modular finger units that each operate independently. Two servo motors actuate each finger by applying tension to four tendons (Figure 2A). MOE has a modular design that allows the fingers to be rearranged into various configurations to suit specific task requirements. In experiments, we use a three-finger variant, inspired by research on object controllability using three-fingered rigid end-effectors [47].
@@
Building on the original MOE finger design, we propose embedding low-cost conductive elastic rubber directly into the silicone elastomer body of each finger. The conductive rubber acts as sensors to measure deformation by varying their electrical resistance as they stretch, providing real-time proprioceptive feedback. Each finger incorporates four of these sensors, compactly positioned between the tendons. A data acquisition (DAQ) circuit and board connected to the sensors can record resistance readings at approximately 400 Hz. By seamlessly integrating the sensors into the elastomer during fabrication, we developed a fully deformable and sensorized finger body. The combined state spaces from all twelve (12) strain sensors provides our estimate $\mathbf{S}$ of the soft robot hand’s true deformation state space $\mathbf{M}$ (Section 3).
4.2 Shape Estimation Model